Shell set, time hist, put to 20 deg and rest.

All OK: SPI-> ~/adopt/measures/adsec_calib/spi/20120527_153751

25 deg: non setta (colpa nella zona open loop della cracca)

30 deg: setta.

All OK: SPI-> ~/adopt/measures/adsec_calib/spi/20120527_160055

Read TSS current on all ok: confirmed good.

Read current on shutdown: see fig below.

30 min time hist + rip + spi: -> ALL ok: adopt/measures/adsec_calib/spi/20120527_180429 + 20120527_180642

Astatic level registration. 1 was found untight, fixed.

Cappello del prete fixed

Neoprene missing on RN fixed

Many slew test done. nothing bad.

-- MarcoXompero - 27 May 2012
Topic revision: r4 - 27 May 2012, MarcoXompero
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback